body executing simple harmonic motion has a velocity of 5m/s at a distance of 5 M and 7m/s at a distance of 4m from the mean position. Find the amplitude and period of oscillation.

Grade:11

2 Answers

Arun
5 years ago
Using V= Wsqrt(A^2 - X^2) Put V= 5,X=5 (get equation 1) V=7, X= 4 (get equation 2)Solve 2 equation in 2 variables get A and w
mani joshi
5 years ago
using equation v=w\sqrt{A^2 + x^2}
5=w\sqrt{A^2 + 5^2}
and
7=w\sqrt{A^2 + 4^2}
divide both the equation
you’ll get A=(5/2*\sqrt{3/2})
put this value of A in ny of the above equation to find w then for time period use 2\pi/w